Circular No. 7277

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V4641 SAGITTARII AND GM SAGITTARII Citing Goranskij (1990, IBVS 3464), we noted on IAUC 7253 that the

star he attributed to GM Sgr is ∼ 1′ north of the position given by Luyten (1927, HCO Bull. No. 852). It has now been established that the recent outbursts identified with GM Sgr (e.g., Downes et al. 1995, A.J. 110, 1824; IAUC 7119, 7120, 7254, 7256, 7257, 7263, 7265, 7267, 7271, 7276) were erroneously attributed to this object, as the result of a misidentification by Kholopov (Goranskij 1978, Astron. Tsirk. No. 1024). N. N. Samus, Sternberg Astronomical Institute, Moscow, informs us that the designation V4641 Sgr has now been given to the apparent early-type cataclysmic variable star, recently identified with SAX J1819.3–2525 = XTE J1819– 254. M. Hazen, Harvard-Smithsonian Center for Astrophysics (CfA), and D. Williams, AAVSO, aided by work in 1959 by B. Welther (CfA), have located a Harvard plate that was taken on 1897 Aug. 3 with the Bruce 0.61-m telescope and evidently used by Luyten. This plate shows both GM Sgr and V4641 Sgr, and G. V. Williams (CfA) provides the following measurement of GM Sgr from a Siding Spring (digitized) survey plate taken on 1987 Apr. 8, indicating that Luyten’s original position was good to within 10′′: α = 18^{h}19^{m}21^{s}.48, δ = −25^{o}25′36′′.0 (equinox 2000.0; 13 USNO-A2.0 reference stars; uncertainty ± 0′′.7). Magnitude estimates of GM Sgr from (blue) Harvard plates (spanning several decades following 1897) by Welther suggest a range from 15.2 (the maximum noted by Luyten) to fainter than 17, indicating that GM Sgr is not a short-period variable. Contributors to, and editors of, other journals should note that what has recently been called GM Sgr (particularly on these Circulars in recent weeks) is in fact V4641 Sgr. Hazen’s research on old Harvard plates regarding GM Sgr is continuing, in collaboration with D. Hoffleit, Yale University Observatory.

V382 VELORUM Photometry of V382 Vel by P. M. Kilmartin at Mt. John Observatory,

Lake Tekapo (cf. IAUC 7238; standard deviations ≤ 0.01 mag): Aug. 25.316 UT, V = 8.28, U − B = −0.89, B − V = −0.43, V − R = +1.04, V − I = +0.32, airmass 2.00; 26.312, 8.34, −0.88, −0.43, +1.03, +0.31, 2.00; Oct. 12.666, 9.05, −1.24, −0.55, +0.60, +0.06, 1.79.

SUPERNOVA 1999ec IN NGC 2207 Corrigendum. On IAUC 7268, last line, for IC 2613. read IC 2163.

1999 October 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green